Concerning SIMPLE INTEREST. I know that I(interest)=P(principal)X R(rate)X T(Time).
How do I solve for P, if I have the rest, 2% for 3 years paid 1080 in interest, how much did I borrow? I cannot figure it out. Help please. Not the answer, just the formula.Thank you, Ted

Answer by nerdybill(7384)

I(interest)=P(principal)X R(rate)X T(Time)
.
Or, I can write as:
I = PRT
Solving for P:
start by dividing both sides by R:
I/R = PT
then, divide both sides by T:
I/(RT) = P (this is your equation)
.
So, given 2% for 3 years paid 1080 in interest
P = I/(RT)
P = 1080/(.02*3)
P = 1080/(.06)
P = $18,000